11. The body with mass m = 1 kg moves without friction according to a trajectory. What will be the velocity of the body at point C? Point A is at a height of 6 m, point B is glued to the plane and point C is at a height of 4.2 m.


1
Expert's answer
2021-04-26T16:59:50-0400
"v^2=2gh\\v^2=2(9.8)(6-4.2)\\v=5.9\frac{m}{s}"